A 29-ft board is cut into two peices, one of which is 13ft longer than the other. How long is each peice?
Answer by checkley71(8403)

X +(X+13)=29
2X+13=29
2X=29-13
2X=16
X=16/2
X=8 ANSWER.
PROOF
8+(8+13)=29
8+21=29
29=29